QUANDALEDINGL3:

Given F(X)=(2x^(2)+5x-12)/(X+4), What Are The Domain And Range?

PuppyBoy932:

Whats this?

Joe348:

If you need help solving equation like this you can always use mathway.com or desmos.com I used both and got this : Domain: all real numbers Range: all real numbers

surjithayer:

\[f(x)=\frac{ 2x^2+5x-12 }{x+4 }\] for domain : \[x+4\neq 0\] \[x \neq -4\] Domain: R-{-4} for range: let f(x)=y \[y=\frac{ 2x^2-5x+12 }{ x+4 }\] \[yx+4y=2x^2-5x+12\] \[2x^2-5x-yx+12-4y=0\] \[2x^2-(5+y)x+(12-4y)=0\] for x to be real disc.>=0 \[b^2-4ac \ge 0\] \[[-(5+y)]^2-4(2)(12-4y)\ge 0\] \[25+10y+y^2-96+32y \ge 0\] \[y^2+42y-71\ge 0\] \[y^2+42y+21^2\ge 71+21^2\] \[(y+21)^2\ge 71+441\] \[(y+21)^2\ge 512\] \[\left| y+21 \right|\ge \sqrt{256 \times 2}\] \[\left| y+21 \right|\ge 16\sqrt{2}\] \[y+21\le -16\sqrt{2}\] \[y \le -21 -16\sqrt{2}\] or \[y+21\ge 16\sqrt{2}\] \[y \ge -21+16 \sqrt{2}\] Range is \[(-\infty,-21-16\sqrt{2}]U[-21+16\sqrt{2},\infty)\]